### v2

# A Study on the Calibration of In-context Learning

Abstract
Modern auto-regressive language models are trained to minimize log loss on broad data by predicting the next token so they are expected to get calibrated answers in next-token prediction tasks. We study this for in-context learning (ICL), a widely used way to adapt frozen large language models (LLMs) via crafting prompts, and investigate the trade-offs between performance and calibration on a wide range of natural language understanding and reasoning tasks. We conduct extensive experiments to show that such trade-offs may get worse as we increase model size, incorporate more ICL examples, and fine-tune models using instruction, dialog, or reinforcement learning from human feedback (RLHF) on carefully curated datasets. Furthermore, we find that common recalibration techniques that are widely effective such as temperature scaling provide limited gains in calibration errors, suggesting that new methods may be required for settings where models are expected to be reliable.
1 Introduction
Language models (LMs) that encompass transformer-based architectures (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) can generate coherent and contextually relevant texts for various use cases. Despite their impressive performance, these models occasionally produce erroneous or overconfident outputs, leading to concerns about their calibration (Dawid, 1982; DeGroot and Fienberg, 1983) which measures how faithful a model’s prediction uncertainty is. Such a problem is pressing as users adapt them using a recent paradigm called in-context learning (Brown et al., 2020) to construct performant predictors, especially for applications in safety-critical domains (Bhatt et al., 2021; Kadavath et al., 2022; Pan et al., 2023).
We provide an in-depth evaluation and analysis of how well these models are calibrated - that is, the alignment between the model’s confidence in its predictions and the actual correctness of those predictions. This token-level calibration assessment will enable us to measure the discrepancy between the model’s perceived and actual performance through a Bayesian uncertainty lens, providing a valuable metric for assessing the model’s accuracy and reliability.
Refer to caption (a) Demonstration of In-context Learning
(b) The accuracy and calibration of LLaMA-30B
Figure 1: The accuracy-calibration trade-off of in-context learning. (a) ICL concerns taking task-specific examples as the prompt to adapt a frozen LLM to predict the answer. (b) Classification accuracy and expected calibration error of ICL. As the number of ICL samples increases, the prediction accuracy improves (Left); at the same time, the calibration gets worse (Right).
We find that LMs including GPT-2 (Radford et al., 2019) and LLaMA (Touvron et al., 2023a) are poorly calibrated and there exists a calibration-accuracy trade-off (Fig.1), i.e. as we increase the amount of in-context samples, the prediction accuracy and calibration error both increase. Crucially, this calibration degradation worsens as the model size increases or when fine-tuning occurs using specialized data, such as curated instructions (Dubois et al., 2023), dialogues (Zheng et al., 2023), or human preference data (Ziegler et al., 2019). Though previous work (Braverman et al., 2020) shows the entropy of each generation step is drifting and can be recalibrated via scaling techniques (Platt et al., 1999) such as temperature scaling (Guo et al., 2017), we show that the miscalibration issue in ICL can not be easily addressed using such well-established recalibration approaches that rely on additional validation data.
Moreover, we study the trade-off in reasoning tasks that involve the generation of explanations (Camburu et al., 2018; Nye et al., 2021; Wei et al., 2022) before the answer, showing that the model can produce confidently wrong answers (using confidence histograms and reliability plots) when prompted with explanations on Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). We carefully design our human assessment to observe that, with the increase in model sizes and the quantity of ICL examples, there is a corresponding rise in the proportion of confidently predicted examples among those incorrectly forecasted. Moreover, we find that a high proportion of wrong predictions are of high confidence and showcase those typical confidently wrong examples of LLMs.
In-context learning has been hypothesized as meta-learning by gradient descent in their forward pass (Von Oswald et al., 2023), which might hopefully yield calibrated predictions (Błasiok et al., 2023) if the models are getting close to local optimality with respect to test loss through meta-optimization. However, the fact that choosing ICL samples from the validation set does not naturally lead to calibrated predictions shows that ICL learns in a fairly different way than SGD. We design controlled experiments to illustrate task learning properties of ICL, showing that when examples in the prompt demonstrate consistent task properties, the learning performance would be improved while the accuracy-calibration trade-off still persists.
2 Related Work
Uncertainty quantification in NLP. Uncertainty quantification in NLP, which often adopts the Bayesian principle to sophisticated methods tailored for neural networks, aims to enhance the reliability of model predictions. This may involve non-trivial designs as directly interpreting language model predictions via probabilities (Kadavath et al., 2022) and linguistic expressions (Lin et al., 2022; Mielke et al., 2022; Zhou et al., 2023b) may inadvertently lead to over-reliance on the model’s uncertainties (Si et al., 2023), thus complicating the establishment of trustworthy common ground between humans and models (Buçinca et al., 2021). Notable recent advancements include employing model confidence as a critical factor in various applications like dialogue generation (Mielke et al., 2022), cascading prediction (Schuster et al., 2021), open-domain QA (Fisch et al., 2020; Angelopoulos et al., 2022), summarization (Laban et al., 2022), language modeling (Schuster et al., 2022), image captioning (Petryk et al., 2023).
Calibration of LLMs. Calibration is a safety property to measure the faithfulness of machine learning models’ uncertainty, especially for error-prone tasks using LLMs. Previous works find that pre-training (Desai and Durrett, 2020) and explanation (Zhang et al., 2020; González et al., 2021) improves calibration. Models can be very poorly calibrated when we prompt LMs (Jiang et al., 2021), while calibration can also depend on model size (Kadavath et al., 2022). (Braverman et al., 2020) assesses the long-term dependencies in a language model’s generations compared to those of the underlying language and finds that entropy drifts as models such as GPT-2 generate text. The intricacy of explanations on complementary team performance poses additional challenges due to the overreliance on explanations of users regardless of their correctness (Bansal et al., 2021). (Mielke et al., 2022) gives a framework for linguistic calibration, a concept that emphasizes the alignment of a model’s expressed confidence or doubt with the actual accuracy of its responses. The process involves annotating generations with <DK>, <LO>, <HI> for confidence levels, then training the confidence-controlled model by appending the control token <DK/LO/HI> at the start of the output, followed by training a calibrator to predict these confidence levels, and finally predicting confidence when generating new examples. (Tian et al., 2023) finds that asking LLMs for their probabilities can be better than using conditional probabilities in a traditional way. (Shih et al., 2023) proposes a simple amortized inference trick for temperature-scaled sampling from LMs and diffusion models. To enhance the estimation of uncertainty in language models, (Kuhn et al., 2023) developed a method that aggregates log probabilities across semantically equivalent outputs. This approach utilizes bidirectional entailment through a model to identify outputs that are semantically similar, thereby refining the uncertainty estimation process. (Cole et al., 2023) identifies the calibration challenge in ambiguous QA and distinguishes uncertainty about the answer (epistemic uncertainty) from uncertainty about the meaning of the question (denotational uncertainty), proposing sampling and self-verification methods. (Kamath et al., 2020) trains a calibrator to identify inputs on which the QA model errs and abstains when it predicts an error is likely. (Zhao et al., 2023) proposes the Pareto optimal learning assessed risk score for calibration and error correction but requires additional training. (Kalai and Vempala, 2023) shows the trade-off between calibration and hallucination but they didn’t study it in an ICL setting and how the predicted answer’s accuracy would impact those two safety aspects.
In-context Learning. Large models such as GPT-3 (Brown et al., 2020) have demonstrated the potential of in-context learning, a method where the model infers the task at hand from the context provided in the input, without requiring explicit retraining or fine-tuning for each new task. Some recent works attempt to understand ICL through meta-learning (Von Oswald et al., 2023), Bayesian inference (Xie et al., 2021), mechanistic interpretability (Olsson et al., 2022), algorithm selection (Bai et al., 2023), synthetic data and simple function classes (Garg et al., 2022; Akyürek et al., 2022). Notably, unlike previous works (Zhao et al., 2021; Han et al., 2023; Fei et al., 2023; Zhou et al., 2023a) that focus on improving task accuracy using the same “calibration” terminology, we study the uncertainty of ICL and measure its trade-off with accuracy.
3 Background
Table 1: Accuracy and Calibration of LLaMA-30B model with three sizes across four text classification datasets and four reasoning datasets
Dataset LLaMA-30B
0-shot 1-shot 4-shot 8-shot
Acc ECE Acc ECE Acc ECE Acc ECE
Text Classification
AGNews 0.383 0.10 0.835 0.416 0.828 0.406 0.856 0.425
TREC 0.651 0.392 0.70 0.442 0.76 0.492 0.777 0.542
CB 0.50 0.143 0.696 0.409 0.821 0.383 0.798 0.359
SST-2 0.537 0.047 0.873 0.367 0.958 0.458 0.958 0.458
DBPdia 0.363 0.287 0.792 0.669 0.83 0.69 0.782 0.646
Reasoning with Scratchpad
Strategy QA 0.452 0.039 0.617 0.047 0.679 0.045 0.678 0.088
Commonsense QA 0.354 0.143 0.613 0.268 0.642 0.279 0.733 0.313
World Tree 0.53 0.253 0.594 0.276 0.655 0.31 0.24 0.230
OpenBook QA 0.388 0.122 0.55 0.15 0.641 0.226 0.66 0.246
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Reliability plots and confidence histograms of LLaMA models on 4-shot learning tasks. Results of different sizes 7B (left), 13B (middle), and 30B (right) are plotted.
Setting. Given a pre-trained language model $\mathcal{P}_{\theta}(w_{t}|w_{<t})$, we seek to adapt it using the prompt $w_{0}=[x_{1},y_{1},x_{2},y_{2},\dots,x_{n-1},y_{n-1},x_{n}]$ to generate a predicted answer $y_{n}=\mathcal{P}_{\theta}(w_{0})$. In the context of reasoning, a popular approach is to hand-craft some explanations/rationales/chain-of-thoughts $e$ in the prompt $w_{0}=[x_{1},e_{1},y_{1},x_{2},e_{2},y_{2},\dots,x_{n-1},e_{n-1},y_{n-1},x_{n}]$ to generate explanation $e_{n}$ and answer $y_{n}$, for the test sample: $\overbrace{w_{1},w_{2},\dots,w_{k}}^{e_{n}},y_{n}=\mathcal{P}_{\theta}(w_{0})$.
We extract token-level answer probabilities of LLMs, [^1] We also normalize the probability $\mathcal{P}_{\theta}\left(y_{n}\mid w_{<n}\right)\in\Delta^{K}$ for classification problems with $K$ choices e.g. for binary classification tasks, we filter and extract probabilities $P(``\textrm{Yes}")$ and $P(``\textrm{No}")$, based on which we calculate the following statistics for studying the confidence and calibration of LMs:
Confidence and feature norm. We record the maximum probability of the answer token as its confidence $\textrm{Conf}=\mathcal{P}_{\theta}(y_{n}|w_{<n})$ and the feature norm $z_{n}$ as the intermediate hidden state before the linear prediction layer.
Entropy rate. We denote the entropy of a token $w_{t}$ at position $t$ as $H(w_{t}|w_{<t})=-\mathbb{E}_{w_{t}\sim\mathcal{P}_{\theta}(\cdot|w_{<t})}[\log\mathcal{P}_{\theta}(w_{t}|w_{<t})]$. We typically measure it based on the answer token via setting $w_{t}=y_{n}$. Note that auto-regressive LLMs are trained via maximizing the negative log-likelihood objective $\mathcal{L}=-\mathbb{E}_{t}[\log\mathcal{P}_{\theta}\left(w_{t}|w_{<t}\right)]$ on massive corpora.
Empirical estimate of the expected calibration error (ECE) In the realm of probabilistic classifiers, calibration is a crucial concept. A classifier, denoted as $\mathcal{P}_{\theta}$ with parameters $\theta$ and operating over $C$ classes, is said to be "canonically calibrated" when, for every probability distribution $p$ over the $C$ classes and for every label $y$, the probability that the label is $y$ given the classifier’s prediction is $p$ matches the component of $p$ corresponding to $y$. This is mathematically represented as:
$\forall p\in\Delta^{C-1},\forall y\in Y:P\left(Y=y\mid\mathcal{P}_{\theta}(X)=p\right)=p_{y}.$
Here, $\Delta^{C-1}$ symbolizes the $(C-1)$-dimensional simplex, which encompasses all potential probability distributions over the $C$ classes.
A simpler calibration criterion is the "top-label calibration." In this case, a classifier is deemed calibrated if, for every top predicted probability $p^{*}$, the probability that the true label belongs to the class with the highest predicted probability, given that this maximum predicted probability is $p^{*}$, equals $p^{*}$. Formally:
$\forall p^{*}\in[0,1]:P\left(Y\in\arg\max p\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right)=p^{*}.$
To gauge the calibration of a model, we adopt Expected Calibration Error (ECE) defined as:
$\mathbb{E}\left[\left|p^{*}-\mathbb{E}\left[Y\in\arg\max\mathcal{P}_{\theta}(X)\mid\max\mathcal{P}_{\theta}(X)=p^{*}\right]\right|\right].$
In real-world applications, this quantity cannot be computed without quantization. So, the ECE is approximated by segmenting predicted confidences into $M$ distinct bins, $B_{1},\ldots,B_{M}$. The approximation is then computed as:
$\widehat{\mathrm{ECE}}=\sum_{m=1}^{M}\frac{\left|B_{m}\right|}{n}\left|\operatorname{acc}\left(B_{m}\right)-\operatorname{conf}\left(B_{m}\right)\right|.$
Here, $\operatorname{acc}\left(B_{m}\right)$ is the accuracy within bin $B_{m}$, and $\operatorname{conf}\left(B_{m}\right)$ is the average confidence of predictions in bin $B_{m}$. The total number of samples is represented by $n$, and the dataset consists of $n$ independent and identically distributed samples, $\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{n}$. In our work, we use this estimator to approximate the ECE.
4 Experimental Results
4.1 Experimental Settings
Models. We study decoder-only autoregressive LMs involving LLaMA (Touvron et al., 2023a), and its variants fine-tuned with instruction, dialog, or RLHF like Alpaca (Dubois et al., 2023), Vicuna (Zheng et al., 2023), and LLaMA2-Chat (Touvron et al., 2023b).
Datasets and tasks. We used both traditional NLU tasks such as AGNews (Zhang et al., 2015), TREC (Voorhees and Tice, 2000), CB (Schick and Schütze, 2021), SST-2 (Socher et al., 2013), DBPedia (Zhang et al., 2015), as well as reasoning question answering tasks like Strategy QA (Geva et al., 2021), Commonsense QA (Talmor et al., 2018), OpenBook QA (Mihaylov et al., 2018), World Tree (Jansen et al., 2018). Notably, the reasoning task performance can be greatly improved in general via prompting methods like scratchpad (Nye et al., 2021; Wei et al., 2022) that enables models to generate natural language explanations before predicting an answer.
In-context learning settings. For $k$-shot learning, we prompt the model via sampling $k$ examples from the training set for each test example. Each experiment is repeated 10 times to reduce variance and we report the mean results. We use $M=10$ bins for calculating calibration errors.
4.2 Numerical Results
The performance of LLaMA. We record the performance and calibration errors for $k$-shot learning ($k=0,1,4,8$), characterizing the calibration-accuracy trade-off in both classic and realistic settings (Tab. 1). Our findings are two-fold: as more in-context examples are included, we observe a concurrent rise in both accuracy and calibration error across most situations, indicating a diminishing reliability of token-level probability in high performance ICL regimes; moreover, when $k=0$ increases to $k=1$, there is a marked boost in both accuracy and calibration error, demonstrating the importance of in-context examples and the significance of the established trade-off. In particular, for reasoning tasks, we explore prompting approaches that explicitly include explanations in reasoning tasks, i.e. scratchpad (Nye et al., 2021) or chain-of-thought (Wei et al., 2022), showing that calibration degrades after generating a long context for reasoning and explaining the final answer.
Table 2: Accuracy and calibration of LLaMA-7B model with different prompt repetition strategies.
Model Strategy 1-shot 4-shot 8-shot AVG ACC AVG ECE
Acc ECE Acc ECE Acc ECE
LLaMA-7B Repeat-context 78.200 0.281 78.370 0.283 77.470 0.274 78.013 0.279
Repeat-prompt 78.200 0.281 89.470 0.395 90.330 0.403 86.000 0.360
Normal 78.200 0.281 95.200 0.452 96.130 0.462 89.843 0.398
The effect of temperature scaling. We experiment with three strategies in applying temperature scaling methods (Guo et al., 2017) to fix miscalibration:
  1. 1.
    We learn one temperature for each $k$-shot ICL, i.e., we learn different temperatures for different shot numbers in ICL;
  2. 2.
    Learn a temperature from the training split (zero-shot) and apply it to all test samples with different shot numbers;
  3. 3.
    For each experiment, we fix the prompt and learn the temperature for the fixed prompt. That is, for every possible ICL prompt, we learn a corresponding temperature for calibration.
Refer to caption (a) Classification accuracy Refer to caption (b) Calibration error
Figure 3: Accuracy and calibration errors of LLaMA and its fine-tuned variants.
Looking into Fig. 6, none of the above strategies achieves satisfactory calibration performance, which is in contrast to the well-studied supervised learning setting where scaling the confidence scores (via temperature scaling) can effectively reduce calibration errors (Guo et al., 2017). The fact that applying a post-processing calibration method, such as temperature scaling, cannot directly resolve the miscalibration issue suggests that ICL might have substantially different properties compared to making predictions via classical supervised learning models.
Refer to caption (a) 00-shot Refer to caption (b) $1$-shot Refer to caption (c) $4$-shot Refer to caption (d) $8$-shot
Figure 4: Illustration of confidence distribution: the number of samples whose confidence is greater than a threshold on Commonsense QA.
Refer to caption (a) 00-shot Refer to caption (b) $1$-shot Refer to caption (c) $4$-shot Refer to caption (d) $8$-shot
Figure 5: The number of wrongly classified examples whose confidence is above a threshold with different numbers of shots on Commonsense QA.
The effect of fine-tuning. We show that vicuna, alpaca and LLaMA2-Chat are all more accurate but less calibrated than their LLaMA counterpart backbones (Fig. 3), the margin is especially large for reasoning tasks and vicuna. Our finding indicates that fine-tuning might significantly degrade calibration, corroborating the evidence reported in GPT-4 (OpenAI, 2023), albeit it can improve the reasoning accuracy dramatically. Our results provide evidence that though fine-tuned on carefully curated datasets can greatly improve question-answering performance, especially for hard tasks like reasoning problems, attention may need to be paid when assessing the calibration of those models’ predictions. We also note that the pattern of zero-shot performance is fairly different for two fine-tuned models, i.e. vicuna, and alpaca.
The effect of prompt formats. In our study, we explore the effects of different prompt strategies using three distinct methods. First, the Repeat-context approach involves constructing prompts as $w_{0}=[x_{1},x_{1},...,x_{1},y_{1}]$, where the context $x_{1}$ is repeated n times, but the label $y_{1}$ is not included in the repetition. Next, the Repeat-prompt strategy shapes the prompt as $w_{0}=[x_{1},y_{1},...,x_{1},y_{1}]$, where both the context $x_{1}$ and the label $y_{1}$ are repeated n times. Finally, the Normal involves constructing the prompt as $w_{0}=[x_{1},y_{1},x_{2},y_{2},...,x_{n-1},y_{n-1},x_{n},y_{n}]$, systematically incorporating distinct context-label pairs.
The findings, as detailed in Tab. 2, reveal certain insights: firstly, integrating labels within prompts significantly decreases uncertainty and enhances reasoning performance. In contrast, simply repeating the context without labels does not lead to better outcomes. Secondly, the diversity of ICL examples in prompt construction notably affects performance, especially for larger language models. Those observations corroborate that ICL is performant when the number of ICL examples is large and they demonstrate consistent task properties. Importantly, the trade-off persists for such controlled scenarios, i.e. improving learning performance in either way would suffer from poorer calibration.
Table 3: Norm of representation, entropy, and confidence of LLaMA-30B model across three text classification datasets.
Dataset LLaMA-30B
Norm Entropy Confidence
0-shot 1-shot 4-shot 8-shot 0-shot 1-shot 4-shot 8-shot 0-shot 1-shot 4-shot 8-shot
AGNews 78.8 92.3 92.1 92.2 3.920 0.650 0.595 0.444 0.214 0.821 0.819 0.865
CB 88.4 91.7 89.2 87.9 3.857 1.266 0.935 0.823 0.193 0.566 0.629 0.577
DBPdia 77.9 89.5 91.0 90.1 4.105 1.438 0.848 0.718 0.078 0.578 0.705 0.671
4.3 Qualitative Results
Figure 6: The comparison of calibration errors before and after applying different recalibration strategies.
Reliability diagram and confidence histogram. A reliability diagram is a graphical tool used to evaluate the calibration of probabilistic predictions of a model across multiple classes; it compares the predicted probabilities of each class against the actual outcomes, with a perfectly calibrated model having its values lie on the diagonal y = x line. A confidence histogram, on the other hand, displays the distribution of the model’s prediction confidences across all classes, showing how often the model predicts certain probabilities.
We showcase that in SST-2 and Open Book QA (4-shot), both ACC and ECE of LLaMA increase as the model size increases (Fig. 2). Moreover, it’s observed that confidence scores tend to concentrate on high values as we enlarge the model size. In particular, the confidence of almost all predictions are above 0.8 for the 30B model.
4.4 Ablation Studies
For case studies, we research how miscalibration can impact the selective classification of LLMs, where models are supposed to abstain from uncertain predictions in high-stakes settings.
Ablation with model sizes. As we enlarge the size of models, they become more confident and accurate (Fig. 2). As a result, the entropy decreases and ECE increases, showing that token-level calibration might have an inverse scaling relationship with model sizes.
Table 4: Qualitative Results of LLaMA on Commonsense and OpenBook
Dataset Case
Reasoning
$\hat{y}$ $p$
Commonsense Question: What would you do if you have curiosity but are blind and paralyzed?
Options: A) hear news, B) analyse, C) go somewhere, D) examine thing, E) see favorite show
Ground Truth: B
7B
If you have curiosity but are blind and paralyzed, you would go somewhere to see things. You cannot hear news, analyse, go somewhere, examine thing or see favorite show.
C 0.5046
13B
If you are blind and paralyzed, you can’t see anything. So, you can’t go somewhere. You can’t examine thing. You can’t see favorite show. You can’t analyse. So, you can hear news.
A 0.8684
30B
If you are blind and paralyzed, you cannot see your favorite show. You cannot hear news as you are blind. You cannot examine thing as you are paralyzed. You can go somewhere to see things.
C 0.9720
Open Book Question: What would you do if you have curiosity but are blind and paralyzed?
Options: A) hear news, B) analyse, C) go somewhere, D) examine thing, E) see favorite show
Ground Truth: B
7B
A girl and her mom have the same hair length.
D 0.6365
13B
A girl and her mom have the same date of birth.
A 0.9782
30B
A girl and her mom have the same genes.
A 0.9831
Examine the hidden state and confidence score. To better understand the miscalibration issue of ICL, we conduct fined-grained experiments to examine ICL properties: we measure the norm of the representation vectors [^2] The representation vector refers to the intermediate output before the linear prediction layer. for different number of shots in ICL. Meanwhile, we also measure the confidence and entropy of the prediction for $y_{n}$, and the results are summarized in Tab. 3. When switching from 0-shot to 1-shot, all three measurements (representation norm, entropy, and confidence) drastically change; on the other hand, when $k$ increases from $1\rightarrow 4\rightarrow 8$, the change of measures would become smoother. Our discovery shows that adding in-context example can substantially impact model behaviors while model behaves relatively similar for various shots once the task is specified ($k\neq 0)$. Meanwhile, more ICL samples lead to smaller entropy and higher confidence in most cases.
Confidence and wrongly classified reasoning examples. To inspect the failure modes of LMs, we randomly sample 100 reasoning examples of LLaMA and plot the distribution of wrongly predicted samples and the confidence scores via thresholding. Similar to previous observations, as model sizes and the number of ICL examples scale up, LMs would generate more confident samples (Fig. 4 (c, d)). We observe “inverse scaling” behaviors where models with larger sizes are more error-prone and tend to generate more confidently wrong samples (Fig. 5).
Examples of hallucinated explanations for highly confident predictions. Next, we showcase in Tab. 4 that models generate both wrong explanations and incorrect predictions with high confidence. We also observe that most of the wrong predictions are highly confident, thus we manually examine the correctness of explanations on commonsense QA, and found its high correlations with predicted answer accuracy, which is the opposite of token-level explainability that tends to get worse when the accuracy improves.
5 Discussion and Concluding Remarks
In our investigation of the token-level calibration of in-context learning in contemporary Large Language Models (LLMs), we delineate the intricate balance between ICL performance and calibration. Our findings underscore the importance of being circumspect in model deployment, as maximizing ICL performance does not invariably translate to improved calibration. As LMs continue to evolve and gain more capabilities such as having long enough context windows that can include the whole training set as in-context examples for some downstream tasks, our result can be pedagogical when users would like to examine their uncertainty through prediction probabilities. Moreover, the work suggests the following future directions:
Understanding the internal mechanism of ICL for calibration.
In this work, we observe that existing scaling recalibration methods cannot fully resolve the miscalibration issues of ICL, so better understanding and mitigation strategies are needed. A potential approach can be leveraging transparency tools and studying whether predictable errors exist during text generation.
Calibration beyond classification regimes.
Our findings indicate that in multi-choice or multi-class classification tasks, even though the calibration of answer tokens may deteriorate in high-performance settings, there may be a positive correlation between accuracy and the correctness of explanations in reasoning tasks. This suggests potential avenues for future research such as exploring strategies such as the use of hedging words to express uncertainty and examining their relationship with predictive performance.
Acknowledgment
We thank Yu Bai, David Childers, Jean-Stanislas Denain for valuable feedback. Kempner Institute computing resources enabled this work.